OpenStudy (anonymous):

Combine like terms. (-3y2 - 7y - 9) - (4y2 + 6y + 9)

OpenStudy (coconutjj):

-3y^2 -7y - 9 - 4y^2 -6y -9

OpenStudy (unklerhaukus):

\[(-3\color{teal}{y^2} - 7\color{brown}{y} - 9) - (4\color{teal}{y^2} + 6\color{brown}{y} + 9)\\ =-3\color{teal}{y^2} - 7\color{brown}{y} - 9 - 4\color{teal}{y^2} - 6\color{brown}{y} - 9\\ =-3\color{teal}{y^2} - 4\color{teal}{y^2} - 7\color{brown}{y} - 6\color{brown}{y} - 9 - 9\\ =(-3- 4)\color{teal}{y^2} +(- 7- 6)\color{brown}{y} +(- 9 - 9)\\ =\]